# A Novel Deep Hybrid Framework with Ensemble-Based Feature Optimization for Robust Real-Time Human Activity Recognition

### v2

@@ Abstract
Human Activity Recognition (HAR) plays a pivotal role in various applications, including smart surveillance, healthcare, assistive technologies, sports analytics, etc. However, HAR systems still face critical challenges, including high computational costs, redundant features, and limited scalability in real-time scenarios. An optimized hybrid deep learning framework is introduced that integrates a customized InceptionV3, an Attention-Augmented AA-LSTM architecture, and a novel ensemble-based feature selection strategy. The proposed framework first extracts spatial descriptors using the customized InceptionV3 model, which captures multilevel contextual patterns, region homogeneity, and fine-grained localization cues. The temporal dependencies across frames are then modeled using AA-LSTMs to effectively encode motion dynamics. Finally, an ensemble-based genetic algorithm with Adaptive Dynamic Fitness Sharing and Attention (ADFSA) is employed to select a compact and optimized feature set by dynamically balancing objectives such as accuracy, redundancy, uniqueness, and complexity reduction. Consequently, the selected feature subsets, which are both diverse and discriminative, enable various lightweight machine learning classifiers to achieve accurate and robust HAR in heterogeneous environments. Experimental results on the robust UCF-YouTube dataset, which presents challenges such as occlusion, cluttered backgrounds, motion dynamics, and poor illumination, demonstrate good performance. The proposed framework achieves 99.65% recognition accuracy, reduces features to as few as 7, and enhances inference time. The lightweight and scalable nature of the HAR system supports real-time deployment on edge devices such as Raspberry Pi, enabling practical applications in intelligent, resource-aware environments, including public safety, assistive technology, and autonomous monitoring systems.
Keywords: Human Activity Recognition, Hybrid Deep Learning, InceptionV3, LSTM Networks, Ensemble Feature Selection, Adaptive Dynamic Fitness Sharing, Real-Time Processing, Edge Computing
@@ 1 Introduction
Recently, Transformer-based models have become popular in human activity recognition. For video data, researchers have combined 3D convolution with attention-based Transformers to better capture both spatial and temporal patterns, while efficiency-focused architectures have enabled lightweight, real-time processing in open-set HAR tasks [9, 10]. In the wearable sensor domain, methods like MoPFormer, SETransformer, and P2LHAP have introduced novel innovations, including tokenization into motion primitives, channel-wise attention, and sequence forecasting, achieving improved accuracy, interpretability, and generalization across datasets [11, 12]. Lightweight, end-to-end Transformers for smartphone IMU data also match or exceed CNN–LSTM baselines with fewer parameters and strong performance [13]. Despite these gains, recent surveys note that Transformer models often entail heavy computation, dependency on large datasets, and limitations when deployed on resource-constrained devices, posing challenges to their broader adoption [14].
To address these limitations, researchers have explored feature fusion and selection strategies for HAR. Promising performance achieved by incorporating deep spatial and motion representations using spatial–temporal attention mechanisms [15] and combining hand-crafted multi-view features with feature selection (FS) based on entropy variability [16]. They often rely on hand-crafted components and produce high-dimensional feature representations, leading to increased computational complexity and limited scalability. Efficient selection of discriminative features is therefore essential to improve inference speed without compromising accuracy.
In this work, we propose a novel hybrid HAR framework that integrates a customized InceptionV3 CNN for spatial feature extraction with an Attention-Augmented (AA-LSTM) network for temporal modeling. The CNN captures discriminative spatial patterns from individual frames, while the AA-LSTM models temporal dependencies across sequences. Designed for real-time video processing, the framework handles challenges such as occlusion, cluttered backgrounds, motion blur, and multiview variations, as exemplified by the UCF YouTube action dataset [17]. Both models are end-to-end trainable and optimized via transfer learning and fine-tuning to ensure efficient inference. Furthermore, to reduce feature redundancy and computational overhead, a genetic algorithm (GA)–based FS strategy is applied, selecting only the most discriminative spatiotemporal features for final classification. By eliminating the need for costly preprocessing steps such as optical flow or skeleton extraction, the proposed framework achieves high accuracy while maintaining practical real-time applicability.
@@
  • A novel ensemble-based GA-driven FS strategy, ADFSA, is employed to achieve compact and discriminative feature subsets. The ADFSA mechanism dynamically reweights multiple objectives: classification accuracy, redundancy minimization, uniqueness enhancement, and complexity reduction through attention-guided adaptation, thereby accelerating convergence and mitigating local stagnation during the evolutionary process.
@@ 2 Related Work
Despite notable progress, existing HAR models still face challenges related to computational efficiency, redundancy in feature representation, and deployment feasibility on real-time or edge platforms. This highlights the need for lightweight yet effective frameworks that can balance recognition accuracy with scalability and efficiency. To overcome these shortcomings, this study proposes a novel deep hybrid framework with ensemble ADFSA FS scheme. The method integrates multiple optimization objectives, including accuracy, redundancy, uniqueness, and complexity, that dynamically adjust their relative importance throughout the evolutionary process. Additionally, by leveraging a two-stage deep feature extraction pipeline (customized InceptionV3 + AA-LSTM), the method effectively captures both spatial and temporal representations. This hybrid design enables the proposed framework to achieve better recognition performance, faster inference, and reduced model complexity, making it well-suited for real-time HAR applications.
@@ 3 Methodology
    Feature Selection and Classification: An ensemble-based genetic algorithm with ADFSA identifies compact, non-redundant, and highly discriminative feature subsets, which are subsequently classified using lightweight machine learning models.
@@ 3.1.1 Frames Extraction from Video Clip
Figure 2: Video to Frames Extraction.
Initially, video data should undergo a frame extraction process, where individual frames are captured at a predetermined rate and resolution, as illustrated in Figure 2, in order to obtain a collection of images with uniform characteristics throughout the dataset.
@@ 3.3 Features Selection
FS plays a critical role in the development of efficient and robust machine learning models, particularly in high-dimensional datasets such as those used in video-based HAR. By identifying and retaining only the most relevant and informative features, FS helps reduce model complexity, minimize overfitting, improve generalization, and significantly lower inference time. It also enhances the interpretability of the model by focusing on features that contribute most to the classification task. The FS process typically involves evaluating the relevance, redundancy, and contribution of each feature using statistical measures, model-based techniques, or optimization algorithms. In this study, a systematic FS approach was applied, which began with extracting a comprehensive set of spatial and temporal features. These features were then evaluated using methods such as PCA [33], RFE [34], and a proposed ensemble-based fitness function to identify the most discriminative subset. The selected features were subsequently used to train lightweight classifiers, ensuring both high accuracy and computational efficiency.
@@ 3.3.1 Dynamic Fitness Sharing for Feature Selection
3.3.2 Proposed Novel Ensemble Adaptive Dynamic Fitness Sharing with Attention Mechanism (ADFSA)
An ensemble fitness function is proposed for feature set optimization in HAR. The proposed technique integrates multiple criteria, classification accuracy, feature redundancy, uniqueness, and complexity into a unified optimization framework using a dynamically adaptive weighting strategy. This ensemble fitness function is designed to guide the give GA-based search process to identify an optimal subset of features that are both discriminative and compact.
@@ 3.3.2 Proposed Novel Ensemble Adaptive Dynamic Fitness Sharing with Attention Mechanism (ADFSA)
Figure 4: Flow of GA incorporated Proposed Ensembled ADFSA
@@ 3.3.3 Mathematical Interpretation of our Proposed ADFSA Technique
To improve convergence and enhance search efficiency, the weights of the ensemble fitness function are dynamically adapted throughout the evolutionary process. The goal is to favor exploration (diversity and redundancy control) in early generations and gradually shift focus toward exploitation (accuracy and compactness) in later generations.
@@
    for each childchild in [$child1$, child2child2do
@@ 3.3.4 Key Advantages of the Proposed Framework
The proposed framework introduces a comprehensive and adaptive FS technique that integrates four fundamental aspects, accuracy, redundancy, uniqueness, and complexity into a unified fitness function. This multi-objective formulation ensures a balanced trade-off between classification performance and model simplicity. One of the key innovations lies in the dynamic adjustment of the weights associated with each fitness component during the evolutionary process. This adaptive mechanism fosters exploration in the initial generations and shifts towards exploitation in later stages, which enhances convergence and helps avoid premature stagnation in local optima. Additionally, the method incorporates a redundancy penalty based on mutual information, effectively minimizing the selection of highly correlated features and ensuring that only the most informative attributes are retained. To further improve population diversity, a uniqueness term is introduced that rewards diverse feature subsets across generations, increasing the likelihood of identifying globally optimal solutions. The inclusion of a complexity component drives the selection of smaller, more efficient feature sets, thereby reducing the computational cost and facilitating faster model inference, an essential requirement for real-time and embedded HAR systems. Empirical evaluations demonstrate that the proposed technique outperforms traditional FS techniques such as PCA and RFE in terms of classification accuracy, even with fewer selected features. Moreover, the approach is classifier-agnostic and compatible with a variety of learning algorithms, including SVM, KNN, Random Forest (RF), and Logistic Regression (LogReg). Its scalability makes it particularly suitable for handling high-dimensional feature vectors derived from deep learning models, addressing the limitations of conventional methods related to scalability and interpretability. Overall, the proposed ensemble-based, dynamically adaptive, and diversity-aware technique presents a novel and effective solution for FS in HAR and other pattern recognition domains.
@@ 4 Experimental Setup
The experimental setup of this research is structured into four main components. First, the dataset is introduced, highlighting its composition and class distribution. Second, a two-phase framework is applied, where spatio-temporal features are extracted using a customized deep learning architecture and then refined through the proposed feature selection technique. Third, the hardware and software environment is outlined to specify the implementation platform and computational resources used. Together, these components provide a clear foundation for evaluating the effectiveness and efficiency of the proposed framework.
@@ 4.1 Dataset
What sets this dataset apart is its variability in camera angles, background clutter, lighting conditions, and motion dynamics, which better reflect real-world scenarios compared to more controlled datasets. Each video sequence is typically labeled with a single dominant activity, and many clips, as showing in Figure 5, contain temporal and spatial variations that test the robustness of recognition models.
Figure 5: YouTube Action Dataset.
@@
Table 2: Configuration of the UCF-YouTube (UCF11) Dataset
@@ 4.2 Two-Phase Feature Extraction and Selection Framework
Overall the research was implemented into two different phases. In the first phase, spatio-temporal features were extracted using a customized Inception-V3 and AA-LSTM network, and in the second phase, the extracted features were passed through the proposed ADFSA FS technique to obtain compact and robust features without affecting the performance of the underlying classifier.
@@ 4.2.1 Spatial Feature Extraction Phase
This work employs a deep CNN architecture to extract discriminative features from video frames. Extensive experiments were carried out by training several deep CNN models, as summarized in Table 3. Among the evaluated architectures, InceptionV3 consistently outperformed the others, demonstrating good recognition accuracy. Trainng of selected networks and the overall proposed framework is presented in Figure 6. Consequently, it was selected and customized for the HAR task, capturing diverse and robust features region and edge, leading to enhanced performance. The training and loss curves of the optimized model are illustrated in Figures 7(a) and 7(b), respectively.
@@ 4.2.3 Features Selection via ADFSA
To enhance the quality of the extracted features, an optimization step was carried out using the proposed ADFSA algorithm. This method leverages a GA guided by an ensemble-based fitness function that integrates four complementary evaluation criteria. First, classification accuracy is assessed using stratified 5-fold cross-validation with a linear SVM, ensuring robust performance estimation. Second, a redundancy penalty based on mutual information is incorporated to discourage the selection of highly correlated features, promoting a more informative subset. Third, a uniqueness component is introduced to enhance population diversity by penalizing the recurrence of identical feature subsets across generations. Finally, a complexity term penalizes larger subsets, thereby encouraging compact and computationally efficient FSs. Notably, the relative weights of these components are adaptively updated throughout the evolutionary process, guided by the normalized progression of the population’s fitness. This dynamic adjustment ensures a balance between exploration and exploitation, ultimately leading to the selection of high-quality, low-redundancy feature sets well-suited for classification tasks. To assess the effectiveness of the proposed technique, two widely used FS techniques were implemented, PCA [25] and RFE [26]. The effectiveness of various selection methods was assessed using a range of classification algorithms, comprising SVM, RF, KNN, and LogReg.
4.3 Hardware and Implementation
Table 4: Configuration details of the experimental setup.
Component Details
Backbone Network Customized inceptionV3 (pretrained)
Temporal Model Customized AA-LSTM with 128 units
Input Size $224\times 224\times 3$
Frames per Video 20
Optimizer Adam ($\text{lr}=1\times 10^{-4}$)
Batch Size 32
Loss Function Categorical Cross-Entropy
Training Epochs 100
Hardware NVIDIA RTX 2080 Ti, 32 GB RAM
5 Results and Discussions
The detailed results are presented in the following subsections. First, 2D feature space visualizations are shown to highlight the separability of the optimized features. Next, evaluation metrics and implementation settings are described. This is followed by a performance summary along with memory and training time analysis. Additional measures such as F1-score, precision, and recall are then reported to provide a broader view of classification performance. An ablation study is included to examine the contribution of each component of the fitness function. Finally, the proposed technique is compared with state-of-the-art approaches to demonstrate its overall effectiveness.

### v3

# A Novel Deep Hybrid Framework with Dynamic Multi-Objective Fitness Optimization for Robust Real-Time Human Activity Recognition

@@ Abstract
Real-time Human Activity Recognition (HAR) has wide-ranging applications in areas such as context-aware environments, public safety, assistive technologies, and autonomous monitoring and surveillance systems. However, existing real-time HAR systems face significant challenges, including limited scalability and high computational costs arising from redundant features. To address these issues, the Inception-V3 model was customized with region-based and boundary-aware operations, using average pooling and max pooling, respectively, to enhance region homogeneity, suppress noise, and capture discriminative local features, while improving robustness through down-sampling. Furthermore, to effectively encode motion dynamics, an Attention-Augmented Long Short-Term Memory (AA-LSTM) network was employed to learn temporal dependencies across video frames. Features are extracted from video dataset and are then optimized through a novel proposed dynamic composite feature selection method called Adaptive Dynamic Fitness Sharing and Attention (ADFSA). This ADFSA mechanism is embedded within a genetic algorithm to select a compact, optimized subset of features by dynamically balancing multiple objectives, accuracy, redundancy reduction, feature uniqueness, and complexity minimization. As a result, the selected subset of diverse and discriminative features enables lightweight machine learning classifiers to achieve accurate and robust HAR in heterogeneous environments. Experimental results demonstrate up to 99.65% accuracy using as few as seven selected features, with improved inference time on the challenging UCF-YouTube dataset, which includes factors such as occlusion, cluttered backgrounds, complex motion dynamics, and poor illumination conditions.
Keywords: Human Activity Recognition, Hybrid Deep Learning, InceptionV3, LSTM Networks, Dynamic Feature Selection, Adaptive Dynamic Fitness Sharing, Real-Time Processing, Edge Computing
@@ 1 Introduction
Recently, Transformer-based models have become popular in human activity recognition. For video data, researchers have combined 3D convolution with attention-based Transformers to capture spatial and temporal patterns, while efficiency-focused architectures have enabled lightweight, real-time processing in open-set HAR tasks [9, 10]. In the wearable sensor domain, methods such as MoPFormer, SETransformer, and P2LHAP have introduced novel innovations, including tokenization in motion primitives, channel-wise attention, and sequence forecasting, achieving improved accuracy, interpretability, and generalization across datasets [11, 12]. Lightweight, end-to-end Transformers for smartphone IMU data also match or exceed CNN–LSTM baselines with fewer parameters and strong performance [13]. Despite these gains, recent surveys note that Transformer models often entail heavy computation, dependency on large datasets, and limitations when deployed on resource-constrained devices, posing challenges to their broader adoption [14].
To address these limitations, researchers have explored feature fusion and selection strategies for HAR. Promising performance achieved by incorporating deep spatial and motion representations using spatial–temporal attention mechanisms [15] and combining hand-crafted multiview features with feature selection (FS) based on entropy variability [16]. They often rely on hand-crafted components and produce high-dimensional feature representations, leading to increased computational complexity and limited scalability. Efficient selection of discriminative features is therefore essential to improve inference speed without compromising accuracy.
In this work, we propose a novel hybrid HAR framework that integrates a customized InceptionV3 CNN for spatial feature extraction with an Attention-Augmented Long-Short-term Memory Network (AA-LSTM) for temporal modeling. CNN captures discriminative spatial patterns from individual frames, while AA-LSTM models temporal dependencies across sequences. Designed for real-time video processing, the framework handles challenges such as occlusion, cluttered backgrounds, motion blur, and multiview variations, as exemplified by the UCF YouTube action dataset [17]. Both models are end-to-end trainable and optimized via transfer learning and fine-tuning to ensure efficient inference. Furthermore, to reduce feature redundancy and computational overhead, a genetic algorithm (GA)–based FS strategy is applied, selecting only the most discriminative spatiotemporal features for final classification. By eliminating the need for costly preprocessing steps such as optical flow or skeleton extraction, the proposed framework achieves high accuracy while maintaining practical real-time applicability.
@@
  • A novel GA-driven FS strategy based on a dynamic composite fitness function, ADFSA, is employed to achieve compact and discriminative feature subsets. The ADFSA mechanism dynamically reweights multiple objectives: classification accuracy, redundancy minimization, uniqueness enhancement, and complexity reduction through attention-guided adaptation, thereby accelerating convergence and mitigating local stagnation during the evolutionary process.
@@ 2 Related Work
Despite notable progress, existing HAR models still face challenges related to computational efficiency, redundancy in feature representation, and deployment feasibility on real-time or edge platforms. This emphasizes the necessity for frameworks that are both lightweight and efficient while maintaining high recognition accuracy and scalability. To address these challenges, this research introduces a new deep hybrid framework that incorporates a dynamic composite ADFSA FS approach. This technique combines various optimization objectives such as accuracy, redundancy, uniqueness, and complexity, which adapt their importance levels during the evolutionary process. Moreover, by utilizing a dual-stage deep feature extraction process (custom InceptionV3 + AA-LSTM), the method effectively captures spatial and temporal features. The hybrid structure of this framework allows for improved recognition performance, quicker inference, and decreased model complexity, making it suitable for real-time Human Activity Recognition (HAR) applications.
@@ 3 Methodology
    Feature Selection and Classification: A dynamic composite fitness function-based genetic algorithm with ADFSA identifies compact, non-redundant, and highly discriminative feature subsets, which are subsequently classified using lightweight machine learning models.
@@ 3.1.1 Frames Extraction from Video Clip
Initially, video data should undergo a frame extraction process, where individual frames are captured at a predetermined rate and resolution, in order to obtain a collection of images with uniform characteristics throughout the dataset.
@@ 3.3 Features Selection
FS plays a critical role in the development of efficient and robust machine learning models, particularly in high-dimensional datasets such as those used in video-based HAR. By identifying and retaining only the most relevant and informative features, FS helps reduce model complexity, minimize overfitting, improve generalization, and significantly lower inference time. It also enhances the interpretability of the model by focusing on features that contribute most to the classification task. The FS process typically involves evaluating the relevance, redundancy, and contribution of each feature using statistical measures, model-based techniques, or optimization algorithms. In this study, a systematic FS approach was applied, which began with extracting a comprehensive set of spatial and temporal features. These features were then evaluated using methods such as PCA [33], RFE [34], and a proposed dynamic composite fitness function to identify the most discriminative subset. The selected features were subsequently used to train lightweight classifiers, ensuring both high accuracy and computational efficiency.
@@ 3.3.1 Dynamic Fitness Sharing for Feature Selection
3.3.2 Proposed Novel composite Adaptive Dynamic Fitness Sharing with Attention Mechanism (ADFSA)
A dynamic composite fitness function is proposed for feature set optimization in HAR. The proposed technique integrates multiple criteria, classification accuracy, feature redundancy, uniqueness, and complexity into a unified optimization framework using a dynamically adaptive weighting strategy. This dynamic composite fitness function is designed to guide the give GA-based search process to identify an optimal subset of features that are both discriminative and compact.
@@ 3.3.2 Proposed Novel composite Adaptive Dynamic Fitness Sharing with Attention Mechanism (ADFSA)
Figure 3: Flow of GA incorporated Proposed dynamic composite ADFSA
@@ 3.3.3 Mathematical Interpretation of our Proposed ADFSA Technique
To improve convergence and enhance search efficiency, the weights of the dynamic composite fitness function are dynamically adapted throughout the evolutionary process. The goal is to favor exploration (diversity and redundancy control) in early generations and gradually shift focus toward exploitation (accuracy and compactness) in later generations.
@@
    for each childchild in [$child1$, $child2$] do
@@ 3.3.4 Key Advantages of the Proposed Framework
The proposed framework introduces a comprehensive and adaptive FS technique that integrates four fundamental aspects, accuracy, redundancy, uniqueness, and complexity into a unified fitness function. This multi-objective formulation ensures a balanced trade-off between classification performance and model simplicity. One of the key innovations lies in the dynamic adjustment of the weights associated with each fitness component during the evolutionary process. This adaptive mechanism fosters exploration in the initial generations and shifts towards exploitation in later stages, which enhances convergence and helps avoid premature stagnation in local optima. Additionally, the method incorporates a redundancy penalty based on mutual information, effectively minimizing the selection of highly correlated features and ensuring that only the most informative attributes are retained. To further improve population diversity, a uniqueness term is introduced that rewards diverse feature subsets across generations, increasing the likelihood of identifying globally optimal solutions. The inclusion of a complexity component drives the selection of smaller, more efficient feature sets, thereby reducing the computational cost and facilitating faster model inference, an essential requirement for real-time and embedded HAR systems. Empirical evaluations demonstrate that the proposed technique outperforms traditional FS techniques such as PCA and RFE in terms of classification accuracy, even with fewer selected features. Moreover, the approach is classifier-agnostic and compatible with a variety of learning algorithms, including SVM, KNN, Random Forest (RF), and Logistic Regression (LogReg). Its scalability makes it particularly suitable for handling high-dimensional feature vectors derived from deep learning models, addressing the limitations of conventional methods related to scalability and interpretability. Overall, the proposed composite, dynamically adaptive, and diversity-aware technique presents a novel and effective solution for FS in HAR and other pattern recognition domains.
@@ 4 Experimental Setup
The experimental setup of this research is structured into four main components as given in the Figure 4. First, the dataset is introduced, highlighting its composition and class distribution. Second, a two-phase framework is applied, where spatio-temporal features are extracted using a customized deep learning architecture and then refined through the proposed feature selection technique. Third, the hardware and software environment is outlined to specify the implementation platform and computational resources used. Together, these components provide a clear foundation for evaluating the effectiveness and efficiency of the proposed framework.
Figure 4: Experimental setup of Proposed Framework
@@ 4.1 Dataset
What sets this dataset apart is its variability in camera angles, background clutter, lighting conditions, and motion dynamics, which better reflect real-world scenarios compared to more controlled datasets. Each video sequence is typically labeled with a single dominant activity, and many clips that test the robustness of recognition models.
@@
Table 2: Configuration of the UCF-YouTube (UCF11) Dataset [11]
@@
4.2 Hardware and Implementation
All algorithms, including feature selection techniques and classifiers such as SVM, RF, KNN, and LogReg, were implemented using Python 3.10 within the Spyder integrated development environment (IDE). The codebase utilized well-established libraries such as TensorFlow/Keras, Scikit-learn, OpenCV, and NumPy. For conventional ML experiments, a workstation with an Intel Core i7 10th-generation processor and 16 GB of RAM was used. In contrast, training of computationally demanding deep models was performed on a high-performance HP desktop, equipped with an Intel Core i9 9th-generation processor, 64 GB of RAM, and an NVIDIA GeForce RTX 2080 Ti GPU.
The deep learning framework consisted of a customized InceptionV3 backbone integrated with an AA-LSTM temporal model. Input frames were resized to $224\times 224\times 3$, which is a widely adopted dimension that provides a balance between computational efficiency and the preservation of discriminative spatial details. Each video clip was represented by 20 frames to capture sufficient temporal information while avoiding excessive redundancy. The Adam optimizer with a learning rate of $1\times 10^{-4}$ was selected due to its robustness in handling sparse gradients and faster convergence in deep architectures. A batch size of 32 was chosen to maintain a trade-off between training stability and GPU memory constraints. Categorical cross-entropy was employed as the loss function since it is well-suited for multi-class classification problems. Finally, the models were trained for 100 epochs, providing enough iterations to achieve convergence without leading to overfitting.
The complete hardware and hyperparameter configuration is summarized in Table 3. This dual-platform and carefully tuned hyperparameter setup ensured efficient execution of lightweight ML tasks as well as effective training of compute-intensive deep learning models.
Table 3: Configuration details of the experimental setup.
Component Details
Backbone Network Customized inceptionV3 (pretrained)
Temporal Model Customized AA-LSTM with 128 units
Input Size $224\times 224\times 3$
Frames per Video 20
Optimizer Adam ($\text{lr}=1\times 10^{-4}$)
Batch Size 32
Loss Function Categorical Cross-Entropy
Training Epochs 100
Hardware NVIDIA RTX 2080 Ti, 32 GB RAM
5 Results and Discussions
The detailed results are presented in the following subsections. First, two-phase features extraction and selection framework is dicussed, and 2D feature space visualizations are shown to highlight the separability of the optimized features. Next, evaluation metrics and implementation settings are described. This is followed by a performance summary along with memory and training time analysis. Additional measures such as F1-score, precision, and recall are then reported to provide a broader view of classification performance. An ablation study is included to examine the contribution of each component of the fitness function. Finally, the proposed technique is compared with state-of-the-art approaches to demonstrate its overall effectiveness.
@@ 5.1 Two-Phase Feature Extraction and Selection Framework
Overall, the research was implemented into two different phases. In the first phase, spatio-temporal features were extracted using a customized Inception-V3 and AA-LSTM network, and in the second phase, the extracted features were passed through the proposed ADFSA FS technique to obtain compact and robust features without affecting the performance of the underlying classifier.
@@ 5.1.1 Spatial Feature Extraction Phase
This work employs a deep CNN architecture to extract discriminative features from video frames. Extensive experiments were carried out by training several deep CNN models, as summarized in Table 4. Among the evaluated architectures, InceptionV3 consistently outperformed the others, demonstrating good recognition accuracy. Trainng of selected networks and the overall proposed framework is presented in Figure 4. Consequently, it was selected and customized for the HAR task, capturing diverse and robust features region and edge, leading to enhanced performance. The training and loss curves of the optimized model are illustrated in Figures 5(a) and 5(b), respectively.
@@ 5.1.3 Features Selection via ADFSA
To enhance the quality of the extracted features, an optimization step was carried out using the proposed ADFSA algorithm. This method leverages a GA guided by a dynamic composite fitness function that integrates four complementary evaluation criteria. First, classification accuracy is assessed using stratified 5-fold cross-validation with a linear SVM, ensuring robust performance estimation. Second, a redundancy penalty based on mutual information is incorporated to discourage the selection of highly correlated features, promoting a more informative subset. Third, a uniqueness component is introduced to enhance population diversity by penalizing the recurrence of identical feature subsets across generations. Finally, a complexity term penalizes larger subsets, thereby encouraging compact and computationally efficient FSs. Notably, the relative weights of these components are adaptively updated throughout the evolutionary process, guided by the normalized progression of the population’s fitness. This dynamic adjustment ensures a balance between exploration and exploitation, ultimately leading to the selection of high-quality, low-redundancy feature sets well-suited for classification tasks. To assess the effectiveness of the proposed technique, two widely used FS techniques were implemented, PCA [25] and RFE [26]. The effectiveness of various selection methods was assessed using a range of classification algorithms, comprising SVM, RF, KNN, and LogReg.